Hello ….. If velocity of sound in a gas depends on its elasticity and density derive the relation for the velocity of sound in a medium by the method of dimension .

Hello …..
      If velocity of sound in a gas depends on its elasticity and density derive the relation for the velocity of sound in a medium by the method of dimension .

Grade:9

1 Answers

Vikas TU
7 years ago
 
As velcoity depends on both that is elasticity it would be:
v = k*a^x d^y
where a is the elasticity and d is thedensity.
Dimensions of
a and d are:
ML^-1T^-2 AND ML^-3T^0 respectively.
 
Dimension velocity => M^0L^1T^-1
Oncomaring we get x+y = 0 and x + 3y = -1
we get
x= ½ and y = -1/2.